The male to female ratio is 5 to 4. There are 6975 students total. How many are female and how many are males?
Answer by checkley79(3341)

5/4=X/(6,975-X) CROSS MULTIPLY
5(6,975-X)=4X
34,875-5X=4X
34,875=4X+5X
9X=34,875
X=34,875/9
X=3,875 MALES.
6,975-3,875=3,100 FEMALES.
PROOF:
5/4=3,875/3,100 DIVIDE.
1.25=1.25